OpenStudy (megan98888):

The polynomial given below has _____ root(s).2x2 + 2x + 4 A.one positive and one negative B.two complex C.two positive D.two negative

OpenStudy (welshfella):

First find the value of the discriminant b^2 - 4ac can you do that/

OpenStudy (megan98888):

No can you help me out more? whats a b and c?

OpenStudy (welshfella):

a b and c sre the coefficients in the equation ax^2 + bx + c = 0 so compare it with 2x^2 + 2x + 4 to get the values of a b and c for this equation.

OpenStudy (welshfella):

compare term by term we see that a = 2 what is the value of b and c?

OpenStudy (megan98888):

Wouldn't it be 2 also?

OpenStudy (welshfella):

which one would be 2? b or c?

jhonyy9 (jhonyy9):

i think may be the first step factorize out the 2

OpenStudy (welshfella):

yes You could do that

jhonyy9 (jhonyy9):

2(x^2 +x +2) =

jhonyy9 (jhonyy9):

and now the discriminant is less than zero so from what result that there are two complex roots

OpenStudy (jdoe0001):

ahemm....just to clarify \(\bf {\color{red}{ 2}} x^2{\color{blue}{ +2}} x{\color{purple}{ +4}}\to \begin{cases} {\color{red}{ a}}\\ {\color{blue}{ b}}\\ {\color{purple}{ c}} \end{cases}\qquad \begin{cases} discriminant \\\hline\\ {\color{blue}{ b}}^2-4{\color{red}{ a}}{\color{purple}{ c}} \end{cases}\)

OpenStudy (welshfella):

I think the poster is confused and has left ..

OpenStudy (jdoe0001):

soooo megan98888... what did you get?

OpenStudy (mathmale):

Given: y = 2x^2 + 2x + 4 compare this to the standard form of a quadratic: y = ax^2 +bx +c Just match up the coefficients. Yes, a=2. What are the values of coefficients b and c?